OpenStudy (anonymous):

Solve for x and y: 1/x+1/y=5 and 5/x-2/y=4

OpenStudy (anonymous):

x=1/2 and y=1/3

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

clearly from the first we have 5/x +5/y =25 now also we have 5/x-2/y =4 subtracting above 2 eqs. we have 7/y =21 or y=1/3 hence putting the value of y in any eqs we have x=1/2